OpenStudy (anonymous):

(3x^ 2a - 4y^ a z ^3a)^ 2 how to solce perfect square

OpenStudy (amistre64):

\[(a+b)^2=a^2+2ab+b^2\]

OpenStudy (anonymous):

is this a formula? but whats a & b

OpenStudy (amistre64):

\[(ab-mn)^2=(ab)^2-2(abmn)+(mn)^2\]

OpenStudy (amistre64):

yes, its a formula the first part is just a single value regardless of how convoluted it appears :) same with the second value

OpenStudy (anonymous):

oh okay. but can you explain to me step by step on how to solve them because am really lost

OpenStudy (amistre64):

(3x^ 2a - 4y^ a z ^3a)^ 2 = (3x^ 2a)^2 - 2(3x^ 2a)(4y^ a z ^3a) +(4y^ a z ^3a)^2

OpenStudy (amistre64):

define your first term (a) and your last term (b) and fill them into the general formula is all it amounts to and dont try to do it all at once, just work on it piece by piece

OpenStudy (amistre64):

you might have been taught a method called FOIL .... thats all it is

OpenStudy (anonymous):

oh okay and when you put the equation into the formula given do you have to put it in lowest terms? or what happens next?

OpenStudy (amistre64):

most likely you have to eventually get all like terms together and neatened up yes

OpenStudy (amistre64):

but i cant really decipher the post well enough to guide you further with it

OpenStudy (anonymous):

oh okay but for this kind of equation how would you set it up to foil it

OpenStudy (amistre64):

take the general equation:\[(a+b)^2\]or if to keep things similar lets go with:\[(a-b)^2\] recall that any number squared is just multiplying itself; 3^2 = 3*3 4^2 = 4*4 x^2 = x*x \[(a-b)^2=(a-b)(a-b)\]